chevrolet raised the base price of its volt by $1400 to $30400

What was the percent increase?
Answer #1

The answer is 2071.4% increase (rounded is 2071.4% increase).

Explanation:

The formula for percent change is:

Percent change = (change/original) * 100

First, you must calculate the change:

30,400 - 1,400 = 29,000

Then, you must put it over the original (the original is 1,400):

Percent change = (29,000/1,400) * 100

Finally, compute the percent change with division and multiplication:--
     
29,000/1,400 = 20.714

20.714 * 100 = 2071.4% increase.